# Scalable Training of Graph Foundation Models for Atomistic Materials Modeling: A Case Study with HydraGNN

### v2

@@ IV-F Full training of best performing HydraGNN models identified by scalable HPO
The 10 best performing HydraGNN models identified by HPO have been fully trained for 40 epochs to reach convergence of the training. We report the trend of the training loss for all 10 models in Fig. 14. The training loss flattens at the end of the training history, indicating that the models have reached their maximum predictive capacity. The trained 10 models will be used to provide ensemble predictions in future tasks with uncertainty measurement.
Figure 14: Full training of 10 selected models from HPO.
V Conclusion and Future Work
@@
In this work we described our approach towards developing and training predictive GFMs by scaling the HydraGNN architecture on hundreds of millions of atomistic materials modeling data using two DOE leadership class supercomputers, viz. NERSC-Perlmutter and OLCF-Frontier. We discussed optimizations and tools used for developing a GFM and running hyperparameter optimization at large scale.
@@
We used distributed data management capabilities to partition large volumes of data across distributed computing resources and efficiently exchange data samples across devices using low-latency communication methods. This helped preserve global data shuffling, which is crucial for maintaining good convergence of the GFM training. By scaling HPO on over 87% of the exascale OLCF-Frontier supercomputer, we have assessed the importance of thoroughly exploring a large set of hyperparameter configurations to identify HydraGNN architectures with high predictive accuracy. Moreover, access to exceptionally performing large scale computing facilities allowed us to develop and test ensemble UQ capabilities to measure the degree of confidence associated with the HydraGNN predictions. Performing HPO and ensemble UQ at unprecedented scale on supercomputing facilities confirms our computational readiness in using HydraGNN to develop trustworthy (i.e, accurate and confident) GFMs to support the US-DOE materials science needs by providing robust and transferable computational capabilities for AI-accelerated materials discovery and design.
@@
Future work will be devoted to deploying the pre-trained GFMs to downstream tasks for fine-tuning, where we will illustrate the efficacy of our GFMs in reducing the amount of training data and computational resources needed to develop robust and transferable DL models for domain-specific applications.
@@ Acknowledgements
Massimiliano Lupo Pasini would like to thank Dr. Vladimir Protopopescu for his valuable feedback in the preparation of the manuscript. This research is sponsored by the Artificial Intelligence Initiative as part of the Laboratory Directed Research and Development (LDRD) Program of Oak Ridge National Laboratory, managed by UT-Battelle, LLC, for the US Department of Energy under contract DE-AC05-00OR22725. This work used resources of the Oak Ridge Leadership Computing Facility, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725, under INCITE award CPH161. This work also used resources of the National Energy Research Scientific Computing Center, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231, under award ERCAP0025216 and ERCAP0027259.

### v3

@@ IV-F Energy-efficient full training of best performing HydraGNN models identified by scalable HPO
The twenty selected HydraGNN models have been fully trained for a maximum of 40 epochs to reach convergence of the training, with the option of early stopping if validation MAE does not decrease across ten consecutive epochs. We report the trend of the training loss for all twenty models in Fig. 23. The training loss flattens at the end of the training history, indicating that the models have reached their maximum predictive capacity. We then used the ensemble of the fully pre-trained twenty GFMs to provide ensemble predictions with epistemic uncertainty measurement, as described in the following Section IV-G.
Figure 22: Energy consumption of each HPO trial plotted against the validation MAE. The black dashed line denotes the Pareto front and the red dots represented the HPO trials selected for ensemble UQ. The HPO trials of the first tier are shown in red and the HPO trials of the second tier are shown in pink.
Figure 23: Full training of twenty selected models from HPO, displaying training and test losses. Nine models are shown for the first tier and eleven models are shown for the second tier.
IV-G Epistemic UQ on ensemble of pre-trained GFMs
Figure 24: Parity plot of ensemble predictions (using twenty pretrained models from Section IV-F) of energy and atomic forces for 20,480 atomistic structures from the test split in pretraining. The red diagonal line represents the perfect model prediction.
Figure 25: Histogram of epistemic uncertainty in energy and force predictions for each dataset used in pretraining.
UQ is essential for the trustworthiness of GFMs. We calculate the uncertainties of the ensemble predictions in pretraining using the twenty models from Section IV-F. Fig. 24 presents the parity plot of ensemble averaged predictions for 20,480 atomistic structures from the test split, showing that our ensemble models have learned the generic information from the diverse pretraining data. While the accuracy is not as high as that reported in task-specific AI work, our pre-trained models are expected to reduce the fine-tuning efforts across a broad set of domain-specific tasks. Moreover, the ensemble averaging further stabilized the predictions of the GFMs.
Figure 25 shows the histogram of epistemic uncertainty of the corresponding energy and force predictions in pretraining, calculated on 4,096 atomistic structure from the test split of each dataset (in total, 20,480 structures). As expected, datasets qm7x and ANI1x, consisting of 9 million organic compounds, exhibit higher uncertainties than the other datasets, which include more than 145 million inorganic compounds. Some of the atomistic structures from the datasets qm7x and ANI1x are associated with significantly inaccurate predictions of atomic forces, which correspond to the data samples sitting on the horizontal point cloud in the parity plot for the prediction of atomic forces. Since the ensemble of GFMs correctly associate high uncertainty to atomistic structures whose predictions of atomic forces are notably inaccurate, this corroborates the fact that our GFMs are self-aware of their own regime of trustworthiness, which is an important indicator of robustness.
This uncertainty information—varying uncertainty values stemming from imbalanced representation in training data—not only provides a confidence measure of our GFM predictions but also guide data collection for further model development in an active learning setting.
V Conclusions and Future Work
@@
In this work we described our approach towards developing and training trustworthy and energy-efficient predictive GFMs by scaling the HydraGNN architecture on over 154 millions of atomistic materials modeling data using two DOE leadership class supercomputers, viz. NERSC-Perlmutter and OLCF-Frontier. We discussed optimizations and tools used for developing a GFM and running HPO at large scale.
@@
We used distributed data management capabilities to partition large volumes of data across distributed computing resources and efficiently exchange data samples across devices using low-latency communication methods. This helped preserve global data shuffling, which is crucial for maintaining good convergence of the GFM training. By scaling HPO on over 91% of the exascale OLCF-Frontier supercomputer, we have assessed the importance of thoroughly exploring a large set of hyperparameter configurations to identify HydraGNN architectures with high predictive accuracy. The use of Omnistat tools at extreme scale allowed us to conduct a thorough analysis to select HydraGNN architectures with high predictive accuracy and low energy cost. Moreover, access to exceptionally performing large scale computing facilities allowed us to develop and test ensemble UQ capabilities to measure the degree of confidence associated with the HydraGNN predictions, which contributes to achieving trustworthiness. Performing HPO and ensemble UQ at unprecedented scale on supercomputing facilities confirms our computational readiness in using HydraGNN to develop trustworthy (i.e, accurate and reliable) and energy-efficient GFMs to support the US-DOE materials science needs by providing robust and transferable computational capabilities for AI-accelerated materials discovery and design.
@@
Future work will be devoted to improving the scaling performance of the GFM pretraining on thousands of GPUs by exploring and developing more sophisticated strategies to mitigate load imbalance between GPUs, while maintaining high quality in the attainable trustworthiness of the pre-trained GFMs. We also plan to apply techniques to read the metadata in parallel to improve the overall bandwidth of reading large-volume of datasets on supercomputers. Once the GFMs are pre-trained, we will deploy them to downstream tasks for fine-tuning, where we will illustrate the efficacy of our GFMs in reducing the amount of training data and computational resources needed to develop robust and transferable DL models for domain-specific applications.
@@ Acknowledgements
Massimiliano Lupo Pasini would like to thank Dr. Vladimir Protopopescu for his valuable feedback in the preparation of the manuscript. This research is sponsored by the Artificial Intelligence Initiative as part of the Laboratory Directed Research and Development (LDRD) Program of Oak Ridge National Laboratory, managed by UT-Battelle, LLC, for the US Department of Energy under contract DE-AC05-00OR22725. This work used resources of the Oak Ridge Leadership Computing Facility, which is supported by the Office of Science of the U.S. Department of Energy, under INCITE award CPH161. This work also used resources of the National Energy Research Scientific Computing Center, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231, under award ERCAP0025216 and ERCAP0027259.